A bag contains 4 red balls, 3 blue balls and 5 yellow balls If a ball is selected at random, find the probability that the ball will be blue or yellow.

a. 0.4167
b. 0.75
c. 0.3333
d. 0.6667
1
Expert's answer
2020-12-28T18:41:53-0500

The total number of balls is 12.

The probability of selecting the blue ball:

"P(blue)=\frac{3}{12}"

The probability of selecting the yellow ball:

"P(yellow)=\frac{5}{12}"

The probability of selecting the blue ball or the yellow ball:

"P(blue)+P(yellow)=\frac{3}{12}+\frac{5}{12}=\\=\frac{8}{12}=\frac{2}{3}=0.6667"


Answer: d 0.6667.